OpenStudy (anonymous):

if 28333=p+23333 then p=?

OpenStudy (unklerhaukus):

Take away 23333 from both sides of the equation

OpenStudy (anonymous):

I got 5 thousand answers are in scientific notation tho

OpenStudy (anonymous):

5 x 10^3 I believe

OpenStudy (anonymous):

10^3=1000, then 1000x5=5000

OpenStudy (unklerhaukus):

good work p=5000 to get this into scientific notation you need to get all the factors of ten together 5000=5*10*10*10=5*10^?

OpenStudy (anonymous):

3 thanks

OpenStudy (unklerhaukus):